OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 − 12x + 1.

OpenStudy (anonymous):

y=1 when x=0 if you think when y=1 again, it should make -12x positive in a way that cancel -2x^2, since -12 = -2*6, -6x must be equal to x^2 so x=-6 Taking the mean between these two numbers gives you -3 for the x coordinate of the vertex because they are simetrical to the center of the parabola. Plug -3 in the equation to find the y coordinate of the vertex and you are done.

OpenStudy (anonymous):

This is based on logic, if you want a more analytical answer just say it.

OpenStudy (anonymous):

woah..okay *starts to read*